# StreamingTalker: Audio-driven 3D Facial Animation with Autoregressive Diffusion Model

### v2

@@ Introduction
Traditional approaches (massaro1212; edwards2016jali) to speech-driven 3D facial animation have predominantly employed deterministic methods that directly map phonemes to the corresponding 3D facial mesh. However, these methods often rely heavily on intermediate representations of phonemes and primarily focus on modeling mouth movements, neglecting other facial dynamics. Moreover, these traditional methods lack the capability for one-to-many generations, which significantly limits their ability to capture the variability and expressiveness of natural facial motion.
In recent years, diffusion models have achieved remarkable success in image generation tasks (ramesh2021zero), leading to their introduction in the domain of speech-driven facial animation (ma2024diffspeaker; sun2024diffposetalk). Notable methods such as FaceDiffuser (stan2023facediffuser) have demonstrated the potential of diffusion processes in this context, achieving fast denoising and good results on short sequences. However, we observe that these designs exhibit a performance drop when generating motions that extend beyond the training horizon. This limitation arises because they are trained on fixed-length sequences, which hinders their ability to generalize to longer and more complex sequences. Furthermore, as these methods employ large denoising networks such as Transformers or U-Nets and need to process the entire sequence before producing results, they incur heavy computation and struggle to support real-time applications.
@@
Existing approaches to speech-driven 3D facial animation can be broadly categorized into linguistics-based and learning-based methods. Linguistics-based methods typically establish a comprehensive rules to govern the animation process. For example, (massaro1212) uses dominance functions to map phonemes to facial movements, while JALI (edwards2016jali) factors mouth movements into lip and jaw animation. However, these methods still rely on intermediate representations of phonemes and primarily focus on mouth movements. In contrast, learning-based methods have emerged to address these limitations. For instance, (taylor2017deep) employs deep neural networks to transform phoneme transcriptions into facial animation parameters, proposing a sliding window approach instead of an RNN. Another notable study (karras2017audio) utilizes convolutional neural networks to animate faces directly from audio data.
Recent studies related to our work have concentrated on training neural networks using cross-modal datasets that combine audio and 3D facial meshes. VOCA (cudeiro2019voca), for example, inputs raw audio and speaker style, represented by subject identity, and utilizes temporal convolutions to animate a static mesh template. MeshTalk (richard2021meshtalk) creates a categorical latent space to distinguish between audio-correlated and uncorrelated facial movements. Both tend to overlook long-term audio context due to their reliance on short audio windows. FaceFormer (fan2022faceformer) addresses this issue by considering long-term audio context with a transformer decoder (vaswani2017attention), enabling it to generate temporally stable animations. In addition, it employs Wav2Vec2.0 (baevski2020wav2vec), a self-supervised pre-trained speech model, which helps mitigate the scarcity of data in existing audio-visual datasets by leveraging large-scale unlabeled speech data to learn rich acoustic and linguistic features. CodeTalker (xing2023codetalker) integrates a temporal autoregressive model with a latent codebook using VQ-VAE (van2017vqvae), inspired by Learning2Listen (ng2022learning2listen). The above-mentioned methods are deterministic models so their diversity is limited since human speech and facial expressions are variable and dynamic.
@@
Diffusion probabilistic models (ho2020denoising; sohl2015diffusion), which differ from previous generative approaches such as GANs (goodfellow2020gan) and VAEs (kingma2013vae), have achieved remarkable results in various generative tasks. These models adopt a Markov chain to gradually add noise to data samples and subsequently use a neural network to approximate the reverse process to denoise the samples. Given their strong abilities in modeling continuous distributions and one-to-many mapping relationships, diffusion models are particularly suitable for animation tasks.
Diffusion models have been applied in related tasks such as 2D talking face generation (du2023dae). FaceDiffuser (stan2023facediffuser) was the first to introduce diffusion models to the domain of speech-driven 3D facial animation. DiffPoseTalk (sun2024diffposetalk) introduces a speaking style encoder and overcomes the limitations of existing diffusion models that cannot be directly transferred to speech-driven expression animation. GLDiTalker (lin2024glditalker) proposes a graph-enhanced quantized space and applies a two-stage training strategy. DiffSpeaker (ma2024diffspeaker) is one of the most recent and effective works in this field, which introduces a novel biased conditional attention mechanism that utilizes encodings to integrate speaking style and diffusion step information. However, these works typically denoise the entire sequence together, which can lead to performance drops when generating samples that extend beyond the training horizon. In contrast, our model innovates using an autoregressive approach to extract conditions from past motions to guide the diffusion process, enabling flexible generation on arbitrary long sequences.
@@
To train generative models more easily, it is common to learn a latent space representation of the raw data (rombach2022). CodeTalker (xing2023codetalker) has shown that VQ-VAE can learn a compact and discrete latent space for facial motions. Inspired by this, we adopt a similar VQ-VAE architecture to encode raw facial motion sequences into latent representations.
@@
We first introduce how we encode inputs. Relative studies (fan2022faceformer; stan2023facediffuser; xing2023codetalker) have shown that self-supervised pre-trained speech model features like Wav2Vec2 (baevski2020wav2vec) and HuBERT (hsu2021hubert) outperform traditional ones such as MFCC. HuBERT was trained on 960 hours of LibriSpeech (panayotov2015librispeech) dataset and (ma2024diffspeaker) using HuBERT achieves previous state-of-the-art performance. Therefore, we choose HuBERT as the audio encoder in this paper and employ a pre-trained hubert-base-ls960 version of it. The audio encoder takes the audio snippet $\mathbf{a}$ as input and outputs a sequence of audio embeddings $E_{a}(\mathbf{a})\in\mathbb{R}^{T\times C}$. In addition, a simple linear projection layer is used to embed the one-hot style vector $s_{k}$ into the latent space as $E_{s}(s_{k})\in\mathbb{R}^{1\times C}$.
@@
As shown in Figure 1 (b), the past motion latents $\mathbf{z}_{past}$ is first passed through a biased causal multi-head (MH) self-attention layer based on ALiBi (press2021train), with a causal mask ensuring that only current or past information is accessible to prevent information leakage. Then, a multi-head cross-attention layer aligns the motion and audio modalities by combining the outputs of the HuBERT encoder and the MH self-attention layer. A notable design is an alignment mask of the decoder (ma2024diffspeaker), which ensures proper alignment of the speech and motion modalities.
@@
In practice, we train a single-layer MLP to approximate this distribution, ensuring real-time generation speed while maintaining accuracy (Equation 5). Particularly, our model directly predicts the clean sample $\mathbf{z}_{0}$, following MDM (tevet2209mdm) and EDGE (tseng2023edge), as it enables us to offer more precise constraints on facial motions. The denoising process is described as follows:
@@
Stage 1.  We pre-train the VQ-VAE model following the approach in (zhang2018mode; razavi2019vqvae2) for 400 iterations. The encoder and decoder of the VQ-VAE are both a six-layer transformer architecture with eight attention heads, with its feature dimension set to 1024 and the feedforward network dimension set to 1536.
In this stage, we supervise the training with a motion reconstruction loss and a quantization loss similar to CodeTalker (xing2023codetalker):
@@
where $\beta=0.25$ is a weighting hyperparameter controlling the update speed of codebook and encoder, and $sg(\cdot)$ stands for a stop-gradient operation which is defined as identity with zero partial derivatives in the backward propagation. Note that the quantization process (Equation 1) is not differentiable, we employ the straight-through gradient estimator following (bengio2013estimating), to copy the gradients from the decoder’s input to the encoder’s output.
Stage 2.  Next, we train the AR condition predictor and diffusion head jointly for 200 iterations while keeping the VQ-VAE motion encoder fixed. The AR condition predictor is a two-layer transformer-based decoder with four attention heads, while the diffusion head is a single-layer MLP sharing the same hidden size. For this phase, we adopt a scaled linear schedule (ho2020denoising) with a total of $N=1000$ steps, and we employ the Denoising Diffusion Implicit Models (DDIM) (song2020ddim), using only 50 steps to sample during inference.
To better improve the quality of latents, we introduce a new latent loss function inspired by MAR (li2024mar) that calculates the L1 distance between the generated latents $\tilde{z}$ and the ground truth latents $\mathbf{z}_{q}$:
@@
We then apply the following geometric losses in the 3D space: the vertex loss $\mathcal{L}_{vert}$ (cudeiro2019voca) for the positions of the mesh vertices and the velocity loss $\mathcal{L}_{vel}$ (cudeiro2019voca) for better temporal consistency:
@@
where $\lambda_{latent}=\lambda_{vert}=\lambda_{vel}=1.0$. All models are trained using the AdamW (loshchilov2017adamw) optimizer with a batch size of 1 and a learning rate of 0.0001 on a single Nvidia RTX 4090 GPU. More details of the network architecture are provided in the Appendix.
@@
We use two publicly available datasets, BIWI (fanelli2010biwi) and VOCASET (cudeiro2019voca), to evaluate the effectiveness our method. Both datasets provide 3D scan-audio pairs of utterances spoken in English.
@@
VOCASET Dataset.  The VOCASET dataset includes 480 paired audio-visual sequences recorded from 12 subjects. The facial motion is captured at 60 frames per second and typically lasts about 4 seconds. Unlike the BIWI dataset, each 3D face mesh in VOCASET is registered to the FLAME (li2017flame) topology, which consists of 5023 vertices.
@@
Following previous works (fan2022faceformer; xing2023codetalker), we adopt two metrics for quantitative evaluation.
Lip Vertex Error (LVE).  We measure lip synchronization following (richard2021meshtalk) by calculating the L2 error between the predicted and ground-truth lip vertices.
Face Dynamics Distance (FDD).  We follow (xing2023codetalker) evaluating the deviation of upper face motions by calculating each upper face vertex’s standard deviation of the element-wise L2 norm along the temporal axis. Smaller FDD indicates that the predicted expressions exhibit high consistency with the natural trends of facial dynamics.
Mouth Open Difference (MOD).  Following (sun2024diffposetalk), we compute the average absolute difference in mouth opening between the predicted and ground truth. A smaller value indicates better alignment in the degree of mouth opening. This helps reflect whether the model generates over-smooth mouth movements.
@@
Baselines.  We compare our method with the following state-of-the-art methods: VOCA (cudeiro2019voca), MeshTalk (richard2021meshtalk), FaceFormer (fan2022faceformer), CodeTalker (ng2022learning2listen), FaceDiffuser (stan2023facediffuser), DiffSpeaker (ma2024diffspeaker), Imitator (thambiraja2023imitator) and DiffPoseTalk (sun2024diffposetalk). Among these, VOCA uses a MLP as decoder while MeshTalk uses a U-Net, both of which use convolutional neural networks to extract features from audio. FaceFormer, CodeTalker and Imitator use transformers to model the long-term context. FaceDiffuser uses diffusion to generate facial animation and applies GRU as the denoiser. DiffSpeaker chooses transformer decoder as the backbone network and outperforms previous methods. DiffPoseTalk predicts FLAME parameters directly. To enable comparison under our vertex-based setting, we adapt its official implementation to output mesh vertices. For all methods, we follow their original training and testing protocols as closely as possible, and apply official pre-trained weights if available. For speakers unseen during training (i.e., not learned in the style embedding space), we generate multiple animations by conditioning on each speaker identity in training set separately, then compute evaluation metrics for each and report the average as the final result.
Table 1: Quantitative comparison of StreamingTalker with baseline methods on two benchmark datasets. For VOCASET, LVE and MOD values are in units of $10^{-5}$$m$. For BIWI, LVE and MOD values are in units of $10^{-4}$$m$, while FDD values are in units of $10^{-5}$$m$. Among them, bold indicates the best performance. $\downarrow$ means lower is better.
@@
($\times 10^{-4}\,\text{mm}$) ($\times 10^{-5}\,\text{mm}$) ($\times 10^{-5}\,\text{mm}$) ($\times 10^{-5}\,\text{mm}$) ($\times 10^{-5}\,\text{mm}$)
@@
We further build a real-time demo leveraging the streaming capability of our model. In this demo, a user speaks to ask a question, and the system first generates a response using a large language model (LLM). The response is then converted to speech via Google TTS (gtts), which is fed into our AR diffusion model to generate facial animation. Thanks to the model’s low-latency design, audio is processed in a streaming manner, with animation frames generated and rendered simultaneously at up to 40 FPS.
@@
Although our proposed approach achieves significant improvements in generating 3D facial animations from speech, it still has some limitations. Firstly, the current dataset consists of relatively short sequences, which are not sufficient to capture the full range of facial expressions a person can exhibit. As a result, our model is unable to extract long-term contextual information, so a larger and more diverse dataset could potentially improve the model’s performance. Secondly, our model currently lacks emotional information, which could be a valuable addition to future research. Incorporating emotional cues, such as EMOCA (danvevcek2022emoca), could enhance the model’s ability to generate more expressive animations and further improve the performance.

### v3

@@ Introduction
Traditional approaches (Massaro et al. 12; Edwards et al. 2016) to speech-driven 3D facial animation have predominantly employed deterministic methods that directly map phonemes to the corresponding 3D facial mesh. However, these methods often rely heavily on intermediate representations of phonemes and primarily focus on modeling mouth movements, neglecting other facial dynamics. Moreover, these traditional methods lack the capability for one-to-many generations, which significantly limits their ability to capture the variability and expressiveness of natural facial motion.
In recent years, diffusion models have achieved remarkable success in image generation tasks (Ramesh et al. 2021), leading to their introduction in the domain of speech-driven facial animation (Ma et al. 2024; Sun et al. 2024). Notable methods such as FaceDiffuser (Stan et al. 2023) have demonstrated the potential of diffusion processes in this context, achieving fast denoising and good results on short sequences. However, we observe that these designs exhibit a performance drop when generating motions that extend beyond the training horizon. This limitation arises because they are trained on fixed-length sequences, which hinders their ability to generalize to longer and more complex sequences. Furthermore, as these methods employ large denoising networks such as Transformers or U-Nets and need to process the entire sequence before producing results, they incur heavy computation and struggle to support real-time applications.
@@
Existing approaches to speech-driven 3D facial animation can be broadly categorized into linguistics-based and learning-based methods. Linguistics-based methods typically establish a comprehensive rules to govern the animation process. For example, (Massaro et al. 12) uses dominance functions to map phonemes to facial movements, while JALI (Edwards et al. 2016) factors mouth movements into lip and jaw animation. However, these methods still rely on intermediate representations of phonemes and primarily focus on mouth movements. In contrast, learning-based methods have emerged to address these limitations. For instance, (Taylor et al. 2017) employs deep neural networks to transform phoneme transcriptions into facial animation parameters, proposing a sliding window approach instead of an RNN. Another notable study (Karras et al. 2017) utilizes convolutional neural networks to animate faces directly from audio data.
Recent studies related to our work have concentrated on training neural networks using cross-modal datasets that combine audio and 3D facial meshes. VOCA (Cudeiro et al. 2019), for example, inputs raw audio and speaker style, represented by subject identity, and utilizes temporal convolutions to animate a static mesh template. MeshTalk (Richard et al. 2021) creates a categorical latent space to distinguish between audio-correlated and uncorrelated facial movements. Both tend to overlook long-term audio context due to their reliance on short audio windows. FaceFormer (Fan et al. 2022) addresses this issue by considering long-term audio context with a transformer decoder (Vaswani 2017), enabling it to generate temporally stable animations. In addition, it employs Wav2Vec2.0 (Baevski et al. 2020), a self-supervised pre-trained speech model, which helps mitigate the scarcity of data in existing audio-visual datasets by leveraging large-scale unlabeled speech data to learn rich acoustic and linguistic features. CodeTalker (Xing et al. 2023) integrates a temporal autoregressive model with a latent codebook using VQ-VAE (Van Den Oord et al. 2017), inspired by Learning2Listen (Ng et al. 2022). The above-mentioned methods are deterministic models so their diversity is limited since human speech and facial expressions are variable and dynamic.
@@
Diffusion probabilistic models (Ho et al. 2020; Sohl-Dickstein et al. 2015), which differ from previous generative approaches such as GANs (Goodfellow et al. 2020) and VAEs (Kingma 2013), have achieved remarkable results in various generative tasks. These models adopt a Markov chain to gradually add noise to data samples and subsequently use a neural network to approximate the reverse process to denoise the samples. Given their strong abilities in modeling continuous distributions and one-to-many mapping relationships, diffusion models are particularly suitable for animation tasks.
Diffusion models have been applied in related tasks such as 2D talking face generation (Du et al. 2023). FaceDiffuser (Stan et al. 2023) was the first to introduce diffusion models to the domain of speech-driven 3D facial animation. DiffPoseTalk (Sun et al. 2024) introduces a speaking style encoder and overcomes the limitations of existing diffusion models that cannot be directly transferred to speech-driven expression animation. GLDiTalker (Lin et al. 2024) proposes a graph-enhanced quantized space and applies a two-stage training strategy. DiffSpeaker (Ma et al. 2024) is one of the most recent and effective works in this field, which introduces a novel biased conditional attention mechanism that utilizes encodings to integrate speaking style and diffusion step information. However, these works typically denoise the entire sequence together, which can lead to performance drops when generating samples that extend beyond the training horizon. In contrast, our model innovates using an autoregressive approach to extract conditions from past motions to guide the diffusion process, enabling flexible generation on arbitrary long sequences.
@@
To train generative models more easily, it is common to learn a latent space representation of the raw data (Rombach et al. 2022). CodeTalker (Xing et al. 2023) has shown that VQ-VAE can learn a compact and discrete latent space for facial motions. Inspired by this, we adopt a similar VQ-VAE architecture to encode raw facial motion sequences into latent representations.
@@
We first introduce how we encode inputs. Relative studies (Fan et al. 2022; Stan et al. 2023; Xing et al. 2023) have shown that self-supervised pre-trained speech model features like Wav2Vec2 (Baevski et al. 2020) and HuBERT (Hsu et al. 2021) outperform traditional ones such as MFCC. HuBERT was trained on 960 hours of LibriSpeech (Panayotov et al. 2015) dataset and (Ma et al. 2024) using HuBERT achieves previous state-of-the-art performance. Therefore, we choose HuBERT as the audio encoder in this paper and employ a pre-trained hubert-base-ls960 version of it. The audio encoder takes the audio snippet $\mathbf{a}$ as input and outputs a sequence of audio embeddings $E_{a}(\mathbf{a})\in\mathbb{R}^{T\times C}$. In addition, a simple linear projection layer is used to embed the one-hot style vector $s_{k}$ into the latent space as $E_{s}(s_{k})\in\mathbb{R}^{1\times C}$.
@@
As shown in Figure 1 (b), the past motion latents $\mathbf{z}_{past}$ is first passed through a biased causal multi-head (MH) self-attention layer based on ALiBi (Press et al. 2021), with a causal mask ensuring that only current or past information is accessible to prevent information leakage. Then, a multi-head cross-attention layer aligns the motion and audio modalities by combining the outputs of the HuBERT encoder and the MH self-attention layer. A notable design is an alignment mask of the decoder (Ma et al. 2024), which ensures proper alignment of the speech and motion modalities.
@@
In practice, we train a single-layer MLP to approximate this distribution, ensuring real-time generation speed while maintaining accuracy (Equation 5). Particularly, our model directly predicts the clean sample $\mathbf{z}_{0}$, following MDM (Tevet et al. 2022) and EDGE (Tseng et al. 2023), as it enables us to offer more precise constraints on facial motions. The denoising process is described as follows:
@@
Stage 1.  We pre-train the VQ-VAE model following the approach in (Zhang et al. 2018; Razavi et al. 2019) for 400 iterations. The encoder and decoder of the VQ-VAE are both a six-layer transformer architecture with eight attention heads, with its feature dimension set to 1024 and the feedforward network dimension set to 1536.
In this stage, we supervise the training with a motion reconstruction loss and a quantization loss similar to CodeTalker (Xing et al. 2023):
@@
where $\beta=0.25$ is a weighting hyperparameter controlling the update speed of codebook and encoder, and $sg(\cdot)$ stands for a stop-gradient operation which is defined as identity with zero partial derivatives in the backward propagation. Note that the quantization process (Equation 1) is not differentiable, we employ the straight-through gradient estimator following (Bengio et al. 2013), to copy the gradients from the decoder’s input to the encoder’s output.
Stage 2.  Next, we train the AR condition predictor and diffusion head jointly for 200 iterations while keeping the VQ-VAE motion encoder fixed. The AR condition predictor is a two-layer transformer-based decoder with four attention heads, while the diffusion head is a single-layer MLP sharing the same hidden size. For this phase, we adopt a scaled linear schedule (Ho et al. 2020) with a total of $N=1000$ steps, and we employ the Denoising Diffusion Implicit Models (DDIM) (Song et al. 2020), using only 50 steps to sample during inference.
To better improve the quality of latents, we introduce a new latent loss function inspired by MAR (Li et al. 2024) that calculates the L1 distance between the generated latents $\tilde{z}$ and the ground truth latents $\mathbf{z}_{q}$:
@@
We then apply the following geometric losses in the 3D space: the vertex loss $\mathcal{L}_{vert}$ (Cudeiro et al. 2019) for the positions of the mesh vertices and the velocity loss $\mathcal{L}_{vel}$ (Cudeiro et al. 2019) for better temporal consistency:
@@
where $\lambda_{latent}=\lambda_{vert}=\lambda_{vel}=1.0$. All models are trained using the AdamW (Loshchilov et al. 2017) optimizer with a batch size of 1 and a learning rate of 0.0001 on a single Nvidia RTX 4090 GPU. More details of the network architecture are provided in the Appendix.
@@
We use two publicly available datasets, BIWI (Fanelli et al. 2010) and VOCASET (Cudeiro et al. 2019), to evaluate the effectiveness our method. Both datasets provide 3D scan-audio pairs of utterances spoken in English.
@@
VOCASET Dataset.  The VOCASET dataset includes 480 paired audio-visual sequences recorded from 12 subjects. The facial motion is captured at 60 frames per second and typically lasts about 4 seconds. Unlike the BIWI dataset, each 3D face mesh in VOCASET is registered to the FLAME (Li et al. 2017) topology, which consists of 5023 vertices.
@@
Following previous works (Fan et al. 2022; Xing et al. 2023), we adopt two metrics for quantitative evaluation.
Lip Vertex Error (LVE).  We measure lip synchronization following (Richard et al. 2021) by calculating the L2 error between the predicted and ground-truth lip vertices.
Face Dynamics Distance (FDD).  We follow (Xing et al. 2023) evaluating the deviation of upper face motions by calculating each upper face vertex’s standard deviation of the element-wise L2 norm along the temporal axis. Smaller FDD indicates that the predicted expressions exhibit high consistency with the natural trends of facial dynamics.
Mouth Open Difference (MOD).  Following (Sun et al. 2024), we compute the average absolute difference in mouth opening between the predicted and ground truth. A smaller value indicates better alignment in the degree of mouth opening. This helps reflect whether the model generates over-smooth mouth movements.
@@
Baselines.  We compare our method with the following state-of-the-art methods: VOCA (Cudeiro et al. 2019), MeshTalk (Richard et al. 2021), FaceFormer (Fan et al. 2022), CodeTalker (Ng et al. 2022), FaceDiffuser (Stan et al. 2023), DiffSpeaker (Ma et al. 2024), Imitator (Thambiraja et al. 2023) and DiffPoseTalk (Sun et al. 2024). Among these, VOCA uses a MLP as decoder while MeshTalk uses a U-Net, both of which use convolutional neural networks to extract features from audio. FaceFormer, CodeTalker and Imitator use transformers to model the long-term context. FaceDiffuser uses diffusion to generate facial animation and applies GRU as the denoiser. DiffSpeaker chooses transformer decoder as the backbone network and outperforms previous methods. DiffPoseTalk predicts FLAME parameters directly. To enable comparison under our vertex-based setting, we adapt its official implementation to output mesh vertices. For all methods, we follow their original training and testing protocols as closely as possible, and apply official pre-trained weights if available. For speakers unseen during training (i.e., not learned in the style embedding space), we generate multiple animations by conditioning on each speaker identity in training set separately, then compute evaluation metrics for each and report the average as the final result.
Table 1: Quantitative comparison of StreamingTalker with baseline methods on two benchmark datasets. For VOCASET, LVE values are in units of $10^{-5}$$m$. For BIWI, LVE values are in units of $10^{-4}$$m$, while FDD values are in units of $10^{-5}$$m$. MOD values are in units of $10^{-3}$$m$ for both datasets. Among them, bold indicates the best performance. $\downarrow$ means lower is better.
@@
($\times 10^{-4}\,\text{mm}$) ($\times 10^{-5}\,\text{mm}$) ($\times 10^{-3}\,\text{mm}$) ($\times 10^{-5}\,\text{mm}$) ($\times 10^{-3}\,\text{mm}$)
@@
We further build a real-time demo leveraging the streaming capability of our model. In this demo, a user speaks to ask a question, and the system first generates a response using a large language model (LLM). The response is then converted to speech via Google TTS (pndurette 2021), which is fed into our AR diffusion model to generate facial animation. Thanks to the model’s low-latency design, audio is processed in a streaming manner, with animation frames generated and rendered simultaneously at up to 40 FPS.
@@
Although our proposed approach achieves significant improvements in generating 3D facial animations from speech, it still has some limitations. Firstly, the current dataset consists of relatively short sequences, which are not sufficient to capture the full range of facial expressions a person can exhibit. As a result, our model is unable to extract long-term contextual information, so a larger and more diverse dataset could potentially improve the model’s performance. Secondly, our model currently lacks emotional information, which could be a valuable addition to future research. Incorporating emotional cues, such as EMOCA (Daněček et al. 2022), could enhance the model’s ability to generate more expressive animations and further improve the performance.